# Extended Galerkin neural network approximation of singular variational problems with error control

### v1

@@ 2.3 Extended neural network architectures
$\Lambda:=\left\{\lambda\in\mathbb{R}^{n_{\lambda}},\;\mathbf{c}_{\lambda}\in\mathbb{R}^{n_{\lambda}}\;:\;||(\lambda,\mathbf{c}_{\lambda})||_{\Lambda}:=||\mathbf{c}_{\lambda}||_{\infty}+||\lambda||_{\infty}\leqslant C\right\}.$
Suppose the mapping $\mathcal{N}:(\Lambda,||\cdot||_{\Lambda})\to(V_{n_{\lambda}}^{\Phi,C},|||\cdot|||)$ defined by
$\mathcal{N}(\lambda,\mathbf{c}_{\lambda})=\sum_{i=1}^{n_{\lambda}}c_{\lambda,i}\Phi(\cdot;\lambda_{i})$
is continuous. Then $V_{n_{\lambda}}^{\Phi,C}$ is compact in $(\mathcal{X},|||\cdot|||)$.
The resulting function $\varphi_{i}^{NN}$ is then the sum of a neural network function $\varphi_{\sigma,i}^{NN}\in V_{\mathbf{n},L}^{\sigma,C}$ and a function $\varphi_{\Phi,i}^{NN}\in V_{n_{\lambda}}^{\Phi,C}$ belonging to the “knowledge-based” space. If the parameters $\lambda$ are unknown in advance, then one possibility is to learn them during the training step in a similar manner to $\mathbf{W}$ and $\mathbf{b}$:
(24) $\lambda\leftarrow\lambda+\nabla_{\lambda}\left[\frac{\langle r(u_{i-1}),v(\cdot;\lambda)\rangle}{|||v(\cdot;\lambda^{)}|||}\right].$
The corresponding coefficients $\mathbf{c}$ and $\mathbf{c}_{\lambda}$ are updated simultaneously by solving the following least squares system
(25) $[\mathbf{c};\mathbf{c}_{\lambda}]=\operatorname*{arg\,min}_{\mathbf{c}\in\mathbb{R}^{n},\mathbf{c}_{\lambda}\in\mathbb{R}^{n_{\lambda}}}||\mathbf{A}[\mathbf{c};\mathbf{c}_{\lambda}]-\mathbf{F}||_{\ell^{2}}$
for an appropriate matrix $\mathbf{A}$ and vector $\mathbf{F}$, and the approximation $u_{i}$ is obtained from the space of extended basis functions $S_{i}^{\Phi}:=\{u_{0}/|||u_{0}|||,\varphi_{\sigma,1}^{NN},\varphi_{\Phi,1}^{NN},\dots,\varphi_{\sigma,i}^{NN},\varphi_{\Phi,i}^{NN}\}$:
(26) $u_{i}\in S_{i}^{\Phi}\;:\;a(u,v)=F(v)\;\;\;\forall v\in S_{i}^{\Phi}.$
The solution may be written as a sum of smooth and singular parts $u_{i}=u_{\sigma,i}+u_{\Phi,i}$, where
(27) $u_{\sigma,i}:=c_{\sigma,0}\frac{u_{0}}{|||u_{0}|||}+\sum_{j=1}^{i}c_{\sigma,j}\varphi_{\sigma,j}^{NN},\;\;\;u_{\Phi,i}:=\sum_{j=1}^{i}c_{\Phi,j}\varphi_{\Phi,j}^{NN}.$
The effectiveness of the extended Galerkin neural network approach depends on the universal approximation properties of the neural networks as in the case of the basic Galerkin neural network approach. The following result, proved in [1] for the basic Galerkin neural network, is directly applicable to problems whose solutions take the form (20). The width $n$ appearing in Proposition 2.5 will depend on both the smooth and singular parts of the solution.
@@
Let $10<\varepsilon<$ 1 be given. Then there exist $n(\varepsilon,u_{\infty}+u_{\Phi}-u_{i-1})\in\mathbb{N}$ and $C(\varepsilon,u_{\infty}+u_{\Psi}-u_{i-1})>0$ such that if $n\geqslant n(\varepsilon,u_{\infty}+u_{\Psi}-u_{i-1})$ and $C\geqslant C(\varepsilon,u_{\infty}+u_{\Psi}-u_{i-1})$, then $\varphi_{i}^{NN}$ given by (3) satisfies $|||\varphi_{i}-\varphi_{i}^{NN}|||\leqslant 2\varepsilon/(1-\varepsilon)$.
By way of contrast with the extended Galerkin neural network approach, the width $n$ depends only on the high regularity part of the solution. Thus, we expect the width of the network required to approximate $\varphi_{i}^{NN}$ to a given tolerance to be smaller than that required in Proposition 2.5 (see e.g. Section 3.1, Example 3.3).
The basis function $\varphi_{i}^{NN}$ satisfies the following approximation property.
Proposition 2.6.
Suppose that the following property holds: For each $\varepsilon>0$, there exists $n_{\lambda}(\varepsilon,\sum_{j=1}^{m}c_{\Psi,j}\Psi(\cdot;\lambda_{i}))\in\mathbb{N}$ and $C(\varepsilon,\sum_{j=1}^{m}c_{\Psi,j}\Psi(\cdot;\lambda_{i}))>0$ and $\tilde{\varphi}_{\Phi}\in\tilde{V}_{n_{\lambda}}^{\Phi}$ such that
(28) $|||\sum_{j=1}^{m}c_{\Psi,j}\Psi(\cdot;\lambda_{i})-\tilde{\varphi}_{\Phi}|||\leqslant\varepsilon.$
Then given $u_{0}\in(\mathcal{X},|||\cdot|||)$ and $L_{i}\in\mathbb{N}$, for each $\varepsilon>0$ there exists $\mathbf{n}^{(i)}(\varepsilon,u_{\infty}-u_{\sigma,i-1})\in\mathbb{N}^{L_{i}}$, $C_{1}(\varepsilon,u_{\infty}-u_{\sigma,i-1})$ as well as $n_{\lambda}^{(i)}(\varepsilon,u_{\Psi}-u_{\Phi,i-1})\in\mathbb{N}$, $C_{2}(\varepsilon,u_{\Psi}-u_{\Phi,i-1})>0$, and $\varphi_{i}^{NN}:=\varphi_{\sigma,i}^{NN}+\varphi_{\Phi,i}^{NN}\in V_{\mathbf{n}^{(i)},L_{i}}^{\sigma,C_{1}}\oplus V_{n_{\lambda}^{(i)}}^{\Phi,C_{2}}$ defined by (23) such that
(29) $|||\varphi_{i}-\varphi_{i}^{NN}|||\leqslant\varepsilon.$
@@
We partition the error by $\varphi_{i}=\varphi_{\sigma,i}+\varphi_{\Phi,i}$, where $\varphi_{\sigma,i}:=(u_{\infty}-u_{\sigma,i-1})/|||u-u_{i-1}|||$ and $\varphi_{\Phi,i}:=(u_{\Psi}-u_{\Phi,i-1})/|||u-u_{i-1}|||$ and first demonstrate that there exists a function $\hat{\varphi}\in V_{\mathbf{n}^{(i)},L_{i}}^{\sigma}\oplus V_{n_{\lambda,i}}^{\Phi}\cap B$ such that $|||\varphi_{i}-\hat{\varphi}|||\leqslant\varepsilon$. Applying the universal approximation theorem for multilayer feedforward networks [10] yields the existence of $\mathbf{n}(\varepsilon,u_{\infty}-u_{\sigma,i-1})\in\mathbb{N}^{L_{i}}$ and $\tilde{\varphi}_{\sigma}\in V_{\mathbf{n}^{(i)},L_{i}}^{\sigma}$ such that $|||\varphi_{\sigma,i}-\tilde{\varphi}_{\sigma}|||<\varepsilon/4$. To approximate $\varphi_{\Phi,i}$, we need only approximate $u_{\Psi}$ and $u_{\Phi,i-1}$. By (28), we can find $\tilde{n}_{\lambda}(\varepsilon,u_{\Psi})\in\mathbb{N}$ and $\tilde{\varphi}_{\Psi}\in\tilde{V}_{\tilde{n}_{\lambda}}^{\Phi}$ such that $|||u_{\Psi}-\tilde{\varphi}_{\Psi}|||<\varepsilon/4$, and since $u_{\Phi,i-1}\in\oplus_{j=1}^{i-1}V_{n_{\lambda}^{(j)}}^{\Phi}$ there exists $\eta\in V_{\sum_{j=1}^{i-1}n_{\lambda}^{(j)}}^{\Phi}$ such that $|||u_{\Phi,i-1}-\eta|||=0$. Setting $\tilde{\varphi}_{\Phi}:=\tilde{\varphi}_{\Psi}+\eta$, we have $|||\varphi_{\Phi,i}-\tilde{\varphi}_{\Phi}|||<\varepsilon/4$. Additionally, setting $\tilde{\varphi}:=\tilde{\varphi}_{\sigma}+\tilde{\varphi}_{\Phi}$, we have $|||\varphi_{i}-\tilde{\varphi}|||\leqslant\varepsilon/2$ and, by the reverse triangle inequality, $|||\tilde{\varphi}|||-1\in(-\varepsilon/2,\varepsilon/2)$.
@@
We next demonstrate that the normalization of $\tilde{\varphi}=\tilde{\varphi}_{\sigma}+\tilde{\varphi}_{\Phi}$, which we shall define by
$\hat{\varphi}:=\frac{\tilde{\varphi}_{\sigma}+\tilde{\varphi}_{\Phi}}{|||\tilde{\varphi}_{\sigma}+\tilde{\varphi}_{\Phi}|||},$
is also an adequate approximation to $\varphi_{i}$. By the triangle inequality, we have
$|||\varphi_{i}-\hat{\varphi}|||$ $\leqslant|||\varphi_{\sigma,i}-\tilde{\varphi}_{\sigma}|||+|||\varphi_{\Phi,i}-\tilde{\varphi}_{\Phi}|||+|||\tilde{\varphi}-\frac{\tilde{\varphi}}{|||\tilde{\varphi}|||}|||$
$\leqslant\varepsilon/2+\frac{1}{|||\tilde{\varphi}|||}|||(|||\tilde{\varphi}|||-1)\tilde{\varphi}|||\leqslant\varepsilon.$
@@
Choose $C_{1}:=||\hat{\theta}||_{\mathcal{NN}}$ and $C_{2}:=||(\hat{\lambda},\mathbf{c}_{\hat{\lambda}})||_{\Lambda}$ where $\hat{\theta}$ and $(\hat{\lambda},\mathbf{c}_{\hat{\lambda}})$ are the parameters corresponding to the realizations $\tilde{\varphi}_{\sigma}/|||\tilde{\varphi}_{\sigma}+\tilde{\varphi}_{\Phi}|||$ and $\tilde{\varphi}_{\Phi}/|||\tilde{\varphi}_{\sigma}+\tilde{\varphi}_{\Phi}|||$, respectively. Additionally, choose $\mathbf{n}^{(i)}$ so that $\mathbf{n}^{(i)}_{j}\geqslant\mathbf{n}_{j}$ for $1\leqslant j\leqslant L_{i}$ and choose $n_{\lambda}^{(i)}\geqslant m+\sum_{j=1}^{i-1}n_{\lambda}^{(j)}$.
@@
It remains to show that $\hat{\varphi}$ is sufficiently close in norm to the maximizer $\varphi_{i}^{NN}$. The argument does not differ from the corresponding proof for the basic Galerkin neural network framework in [1]. For completeness, we summarize the argument here. Since $\varphi_{i}^{NN}$ is the maximizer of $\langle r(u_{i-1}),\cdot\rangle=a(u-u_{i-1},\cdot)$, we have $a(\varphi_{i},\varphi_{i}^{NN})\geqslant a(\varphi_{i},v)$ for all $v\in V_{\mathbf{n}^{(i)},L_{i}}^{\sigma,C_{1}}\oplus V_{n_{\lambda}^{(i)}}^{\Phi,C_{2}}\cap B$ and
$|||\varphi_{i}-\varphi_{i}^{NN}|||^{2}$ $=|||\varphi_{i}|||^{2}-a(\varphi_{i},\varphi_{i}^{NN})+|||\varphi_{i}^{NN}|||^{2}$
$\leqslant|||\varphi_{i}|||^{2}-a(\varphi_{i},\hat{\varphi})+|||\hat{\varphi}|||^{2}=|||\varphi_{i}-\hat{\varphi}|||^{2}\leqslant\varepsilon^{2}.$
We obtain the following result demonstrating the convergence of the extended Galerkin neural network.
@@
Moreover, if $\varepsilon<1$, then each $\varphi_{i}^{NN}$ satisfies11footnotemark: 1
@@
$\left|\langle r(u_{i-1}),\varphi_{i}^{NN}-|||u-u_{i-1}|||\right|\leqslant|||u-u_{i-1}|||\cdot|||\varphi_{i}^{NN}-\varphi_{i}|||=\varepsilon|||u-u_{i-1}|||$

### v2

@@ 2.3 Extended neural network architectures
$\Lambda:=\left\{\mu\in\mathcal{I}_{\Phi}^{m},\;\mathbf{d}\in\mathbb{R}^{m}\;:\;||(\mu,\mathbf{d})||_{\Lambda}:=||\mathbf{d}||_{\infty}+||\mu||_{\infty}\leqslant C\right\}.$
Suppose the mapping $\mathcal{N}:(\Lambda,||\cdot||_{\Lambda})\to(V_{m}^{\Phi,C},|||\cdot|||)$ defined by
$\mathcal{N}(\mu,\mathbf{d})=\sum_{i=1}^{m}d_{i}\Phi(\cdot;\mu_{i})$
is continuous. Then $V_{m}^{\Phi,C}$ is compact in $(\mathcal{X},|||\cdot|||)$.
The set $V_{\mathcal{M}}^{\Psi,C_{2}}(\lambda)$ is always compact in $\mathcal{X}$ since the corresponding mapping $\mathcal{N}$ in Lemma 2.4 is always continuous whenever the nonlinear parameter $\mu$ is fixed, as is the case for the space $V_{\mathcal{M}}^{\Psi,C_{2}}$. Thus, the maximization problem in (29) is well-defined. We postpone discussions of compactness of the more complicated set $V_{m}^{\Phi}$ until its use in Proposition 2.11 (see Remark 2.10).
The effectiveness of the extended Galerkin neural network approach depends on the universal approximation properties of neural networks. The following result, proved in [1] for the basic Galerkin neural network in the case of single hidden layer networks, is directly applicable to problems whose solutions take the form (28).
Proposition 2.5 ([1, Proposition 2.5]).
Let $10<\varepsilon<$ 1 be given and consider networks consisting of a single hidden layer, i.e. $\mathbf{n}^{(i)}=n$. Then there exist $n(\varepsilon,u_{\infty}+u_{\Psi}-u_{i-1})\in\mathbb{N}$ and $C(\varepsilon,u_{\infty}+u_{\Psi}-u_{i-1})>0$ such that if $n\geqslant n(\varepsilon,u_{\infty}+u_{\Psi}-u_{i-1})$ and $C\geqslant C(\varepsilon,u_{\infty}+u_{\Psi}-u_{i-1})$, then $\varphi_{i}^{NN}$ given by (3) satisfies $|||\varphi_{i}-\varphi_{i}^{NN}|||\leqslant 2\varepsilon/(1-\varepsilon)$.
It will not have escaped the reader’s attention that the width $n$ appearing in Proposition 2.5 depends on both the smooth and singular parts of the solution.
By way of contrast, the next result, Proposition 2.6 applies to multilayer networks and shows that if the extended Galerkin neural network approach is used, then the width $\mathbf{n}$ now only depends on the smooth part of the solution, whereas if the knowledge-based functions are omitted, then the width $\mathbf{n}$ would also depend on the singular part of the solution (which we expect to be larger). This expectation is borne out in Example 3.3.
@@
Suppose $\varepsilon>0$. Then given $u_{0}\in(\mathcal{X},|||\cdot|||)$ and $L_{i}\in\mathbb{N}$, there exists $\mathbf{n}^{(i)}\in\mathbb{N}^{L_{i}}$ and $C_{1}>0$ depending only on $\varepsilon$ and the error $u_{\infty}-u_{\sigma,i-1}$ in the approximation of the high-regularity solution component as well as $C_{2}>0$ depending only on $\varepsilon$ and the error $u_{\Psi}-u_{\Psi,i-1}$ in the approximation of the low-regularity solution component such that
(33) $|||\varphi_{i}-\varphi_{i}^{NN}|||\leqslant\varepsilon,$
where $\varphi_{i}^{NN}:=\varphi_{\sigma,i}^{NN}+\varphi_{\Psi,i}^{NN}\in V_{\mathbf{n}^{(i)},L_{i}}^{\sigma,C_{1}}\oplus V_{\mathcal{M}}^{\Psi}(\lambda)$ is defined by (29).
@@
We partition the normalized approximation error $\varphi_{i}=(u-u_{i-1})/|||u-u_{i-1}|||=\varphi_{\sigma,i}+\tilde{\varphi}_{\Psi}$, where $\varphi_{\sigma,i}:=(u_{\infty}-u_{\sigma,i-1})/|||u-u_{i-1}|||$ and $\varphi_{\Psi}:=(u_{\Psi}-u_{\Psi,i-1})/|||u-u_{i-1}|||\in V_{\mathcal{M}}^{\Psi}(\lambda)$.
Thanks to [10, 11], we may approximate $\varphi_{\sigma,i}$ by $\tilde{\varphi}_{\sigma}\in V_{\mathbf{n},L_{i}}^{\sigma}$ such that
(34) $|||\varphi_{\sigma,i}-\tilde{\varphi}_{\sigma}|||\leqslant\varepsilon/2$
provided that $\mathbf{n}$ is sufficiently large depending on $\varepsilon$ and $u_{\infty}-u_{\sigma,i-1}$. Thus, setting $\tilde{\varphi}:=\tilde{\varphi}_{\sigma}+\varphi_{\Psi}$, we have the estimate
(35) $|||\varphi_{i}-\tilde{\varphi}|||\leqslant\varepsilon/2.$
Moreover, since $|||\varphi_{i}|||=1$, we have $|||\tilde{\varphi}|||-1\in(-\varepsilon/2,\varepsilon/2)$.
@@
We next demonstrate that the normalization of $\tilde{\varphi}$, which we shall define by $\hat{\varphi}:=\tilde{\varphi}/|||\tilde{\varphi}|||$, is also an adequate approximation to $\varphi_{i}$. By the triangle inequality, we have
$|||\varphi_{i}-\hat{\varphi}|||$ $\leqslant|||\varphi_{i}-\tilde{\varphi}|||+|||\tilde{\varphi}-\frac{\tilde{\varphi}}{|||\tilde{\varphi}|||}|||$
(36) $\leqslant\varepsilon/2+\frac{1}{|||\tilde{\varphi}|||}|||(|||\tilde{\varphi}|||-1)\tilde{\varphi}|||\leqslant\varepsilon.$
@@
Choose $\mathbf{n}^{(i)}$ so that $\mathbf{n}^{(i)}_{j}\geqslant\mathbf{n}_{j}$ for $1\leqslant j\leqslant L_{i}$, $C_{1}\geqslant||(\mathbf{\hat{W}},\mathbf{\hat{b}},\mathbf{\hat{c}})||_{\mathcal{NN}}$, and $C_{2}\geqslant||(\hat{\mu},\mathbf{\hat{d}})||_{\Lambda}$ where $(\mathbf{\hat{W}},\mathbf{\hat{b}},\mathbf{\hat{c}})$ and $(\hat{\mu},\mathbf{\hat{d}})$ are the parameters corresponding to the realizations $\tilde{\varphi}_{\sigma}/|||\tilde{\varphi}|||$ and $\varphi_{\Psi}$, respectively.
@@
It remains to show that the maximizer $\varphi_{i}^{NN}$ is sufficiently close in norm $\hat{\varphi}$ and hence, $\varphi_{i}$. The argument does not differ from the corresponding proof for the basic Galerkin neural network framework in [1]. For completeness, we summarize the argument here. Since $\varphi_{i}^{NN}$ is the maximizer of $\langle r(u_{i-1}),\cdot\rangle=a(u-u_{i-1},\cdot)$, we have $a(\varphi_{i},\varphi_{i}^{NN})\geqslant a(\varphi_{i},v)$ for all $v\in V_{\mathbf{n}^{(i)},L_{i}}^{\sigma,C_{1}}\oplus V_{\mathcal{M}}^{\Psi,C_{2}}(\lambda)$ with $|||v|||=1$, and
$|||\varphi_{i}-\varphi_{i}^{NN}|||^{2}$ $=|||\varphi_{i}|||^{2}-2a(\varphi_{i},\varphi_{i}^{NN})+|||\varphi_{i}^{NN}|||^{2}$
$\leqslant|||\varphi_{i}|||^{2}-2a(\varphi_{i},\hat{\varphi})+|||\hat{\varphi}|||^{2}=|||\varphi_{i}-\hat{\varphi}|||^{2}\leqslant\varepsilon^{2}.$
At first sight, one might expect that the knowledge-based functions need only be utilized once and for all in the first iteration, i.e. $i=1$. However, as the Galerkin iteration proceeds and the smooth part of the approximation encapsulated by the feedforward neural network becomes more refined, it may be necessary to adjust the linear coefficients $\mathbf{d}^{(i)}$ of the knowledge-based functions accordingly. For this reason, the singular functions are incorporated at every subsequent Galerkin iteration.
Given Proposition 2.6, we may also obtain approximation results for $u_{i}\approx u$ and the a posteriori estimator $\langle r(u_{i-1}),\varphi_{i}^{NN}\rangle\approx|||u-u_{i-1}|||$.
@@
Moreover, if $\varepsilon<1$, then each $\varphi_{i}^{NN}$ satisfies
@@
$\left|\langle r(u_{i-1}),\varphi_{i}^{NN}\rangle-|||u-u_{i-1}|||\right|\leqslant|||u-u_{i-1}|||\cdot|||\varphi_{i}^{NN}-\varphi_{i}|||=\varepsilon|||u-u_{i-1}|||$
@@
In some cases, one might have knowledge of the structure of $\Psi$ while not having access to the exact values of $\lambda$. The parameters $\mu^{(i)}$ corresponding to $\varphi_{\Psi,i}^{NN}$ may then be treated as trainable nonlinear parameters and learned during the training step. One might also not know the structure of $\Psi$ in its entirety but instead have a suitable function $\Phi(\cdot;\lambda)$ which approximates $\Psi(\cdot,\lambda)$. In this case, a result analogous to Proposition 2.6 may still be obtained. In what follows, we shall assume that $\Phi(\cdot,\mu)\in\mathcal{X}$, $\mu\in\mathcal{I}_{\Phi}$ is a family of functions such that for each $\varepsilon>0$, there exists $m_{*}(\varepsilon,u_{\Psi})\in\mathbb{N}$ and $\tilde{\varphi}_{\Phi}\in V_{m_{*}}^{\Phi}$ such that
(39) $|||u_{\Psi}-\tilde{\varphi}_{\Phi}|||\leqslant\varepsilon/2.$
We propose seeking $\varphi_{\Psi,i}^{NN}\in V_{m_{*}}^{\Phi,C_{2}}\oplus(V_{m_{*}}^{\Phi,C_{2}}(\mu^{(i-1)})\oplus\dots\oplus V_{m_{*}}^{\Phi,C_{2}}(\mu^{(0)}))$, where $\mu^{(1)},\dots,\mu^{(i-1)}$ are the learned nonlinear hyperparameters corresponding to $\varphi_{\Psi,1}^{NN},\dots,\varphi_{\Psi,i-1}^{NN}$ from the prior Galerkin iterations. The basis function $\varphi_{i}^{NN}$ is then obtained according to
(40) $\varphi_{i}^{NN}=\operatorname*{arg\,max}_{\begin{subarray}{c}v\in V_{\mathbf{n}^{(i)}}^{\sigma,C_{1}}\oplus\left(V_{m_{*}}^{\Phi,C_{2}}\oplus V_{m_{*}}^{\Phi,C_{2}}(\mu^{(i-1)})\oplus\dots\oplus V_{m_{*}}^{\Phi,C_{2}}(\mu^{(0)})\right)\\ |||v|||=1\end{subarray}}\langle r(u_{i-1}),v\rangle.$
Note that the set functions $V_{m_{*}}^{\Phi,C_{2}}\oplus(V_{m_{*}}^{\Phi,C_{2}}(\mu^{(i-1)})\oplus\dots\oplus V_{m_{*}}^{\Phi,C_{2}}(\mu^{(0)}))$ from which the knowledge-based part of the basis function is obtained has width $i\cdot m_{*}$. This is in contrast to (29) where the width of the set of knowledge-based functions was fixed and did not depend on $i$. The necessity of this growth is due to the nonlinear nature of the set $V_{m_{*}}^{\Phi,C}$. In particular, to correct the linear coefficients $\mathbf{d}^{(i)}$ corresponding to the singular part of the previous basis functions as the solution is refined, the knowledge-based functions $\Phi(\cdot;\mu^{(i)})$ must be utilized at every Galerkin iteration. The following result demonstrates that it is possible to obtain an analogous approximation result to Proposition 2.6 when using approximate knowledge-based functions and inexact singular hyperparameters.
Remark 2.10.
Compactness of the set $V_{m}^{\Phi,C}$ depends on the precise structure of $\Phi$ and cannot be proven for general $\Phi$. We briefly note, however, that the special choice $\Phi(x;\mu)=r(x)^{\mu}\sin(\mu\theta(x))$, where $(r,\theta)$ are the standard polar coordinate, results in $V_{m}^{\Phi,C}$ being a compact subset of $\mathcal{X}$. This function commonly arises in the solution of PDEs with singular behavior in non-convex corners. It is not difficult to show that $||\Phi(\cdot;\mu_{1})-\Phi(\cdot;\mu_{2})||_{H^{s}(\Omega)}\to 0$ as $\mu_{1}\to\mu_{2}$ provided that $\mu_{1},\mu_{2}>0$ are chosen appropriately.
Proposition 2.11.
Suppose $\varepsilon>0$ and there exists $\tilde{\varphi}_{\Phi}\in V_{m_{*}}^{\Phi}$ for some sufficiently large $m_{*}\in\mathbb{N}$ such that (39) holds. Then given $u_{0}\in(\mathcal{X},|||\cdot|||)$ and $L_{i}\in\mathbb{N}$, there exists $\mathbf{n}^{(i)}\in\mathbb{N}^{L_{i}}$ and $C_{1}>0$ depending only on $\varepsilon$ and the error $u_{\infty}-u_{\sigma,i-1}$ in the approximation of the high-regularity component as well as $C_{2}>0$ depending only on $\varepsilon$ and the error $u_{\Psi}-u_{\Psi,i-1}$ low-regularity component such that
(41) $|||\varphi_{i}-\varphi_{i}^{NN}|||\leqslant 2\varepsilon,$
where $\varphi_{i}^{NN}$ is defined by (40).
Proof 2.12.
The two key differences between this scenario and the one presented in Proposition 2.6 are (1) the set $V_{m}^{\Phi}$ from which we seek the basis functions is nonlinear since $\lambda$ is a trainable parameter, meaning that the widths $m^{(i)}$ must grow with each Galerkin iteration in order to allow for adjustments to the linear coefficients $\mathbf{d}^{(i)}$ as the smooth part of the approximation is refined (2) the approximation error includes additional error introduced by the inexact function $\Phi$.
We need only prove the existence of $\mathbf{n}^{(i)}$, $C_{1}$, $C_{2}$, and $\hat{\varphi}\in V_{\mathbf{n}^{(i)},L_{i}}^{\sigma,C_{1}}\oplus V_{m_{*}}^{\Phi,C_{2}}\oplus V_{m_{*}}^{\Phi,C_{2}}(\mu^{(i-1)})\oplus\dots\oplus V_{m_{*}}^{\Phi,C_{2}}(\mu^{(0)}))$ with $|||\hat{\varphi}|||=1$ such that $|||\hat{\varphi}-\varphi_{i}|||\leqslant 2\varepsilon$. The remainder of the proof does not differ from that of Proposition 2.6. We again set $\hat{\varphi}=\tilde{\varphi}/|||\tilde{\varphi}|||$ where $\tilde{\varphi}:=\tilde{\varphi}_{\sigma}+\tilde{\varphi}_{\Psi}$, $\tilde{\varphi}_{\sigma}\in V_{\mathbf{n},L_{i}}^{\sigma}$ is defined the same way as in the proof of Proposition 2.6, and the existence of $\tilde{\varphi}_{\Psi}$ is established as follows. Recall that $\varphi_{\Psi}\propto u_{\Psi}-u_{\Psi,i-1}$. Due to (39), a function from the set $V_{m_{*}}^{\Phi}$ is sufficient to approximate $u_{\Psi}$ to within $\varepsilon/2$. Likewise, due to (32) we have $u_{\Psi,i-1}\in V_{m_{*}}^{\Phi}(\mu^{(0)})\oplus\dots\oplus V_{m_{*}}^{\Phi}(\mu^{(i-1)})$. Thus, there exists $\tilde{\varphi}_{\Psi}\in V_{m_{*}}^{\Phi}\oplus(V_{m_{*}}^{\Phi}(\mu^{(i-1)})\oplus\dots\oplus V_{m_{*}}^{\Phi}(\mu^{(0)}))$ such that
(42) $|||\varphi_{\Psi}-\tilde{\varphi}_{\Psi}|||\leqslant\varepsilon/2$
and we have
[…]
We choose $\mathbf{n}^{(i)}$ such that $\mathbf{n}^{(i)}_{j}\geqslant\mathbf{n}_{j}$, $1\leqslant j\leqslant L_{i}$; $C_{1}$ such that $C_{1}\geqslant||(\mathbf{\hat{W}},\mathbf{\hat{b}},\mathbf{\hat{c}})||_{\mathcal{NN}}$; and $C_{2}$ such that $C_{2}\geqslant||(\hat{\mu},\mathbf{\hat{d}})||_{\Lambda}$, where $(\mathbf{\hat{W}},\mathbf{\hat{b}},\mathbf{\hat{c}})$ and $(\hat{\mu},\mathbf{\hat{d}})$ are the parameters corresponding to the realizations $\tilde{\varphi}_{\sigma}/|||\tilde{\varphi}|||$ and $\tilde{\varphi}_{\Psi}/|||\tilde{\varphi}|||$, respectively.
An analogous result to Corollary 2.8 may be derived for the case described by (40). The proof does not differ from that of Corollary 2.8.
Finally, we briefly discuss the training process for computing the maximizer in (40). The nonlinear hyperparameters corresponding to the knowledge-based part of the basis function, $\varphi_{\Psi,i}^{NN}$, are given by $[\mu^{(i)};\mu^{(i-1)};\dots;\mu^{(0)}]$, where $\mu^{(i)}$ corresponding to the part of $\varphi_{\Psi,i}^{NN}$ in $V_{m_{*}}^{\Phi,C_{2}}$. Only $\mu^{(i)}$ should be trained since $\mu^{(i-1)},\dots,\mu^{(0)}$ are fixed according to the previous Galerkin iterations. We update $\mu^{(i)}$ using a gradient-based optimizer; for convenience, we show a gradient descent update here:
(44) $\mu^{(i)}\leftarrow\mu^{(i)}+\nabla_{\mu^{(i)}}\left[\frac{\langle r(u_{i-1}),v(\cdot;\mu^{(i)})\rangle}{|||v(\cdot;\mu^{(i)})|||}\right].$
The corresponding linear coefficients $\mathbf{c}$ and $\mathbf{d}$ are updated simultaneously by solving the following least squares system
(45) $[\mathbf{c};\mathbf{d}]=\operatorname*{arg\,min}_{\mathbf{c}\in\mathbb{R}^{n},\mathbf{d}\in\mathbb{R}^{i\cdot m_{*}}}||\mathbf{A}[\mathbf{c};\mathbf{d}]-\mathbf{F}||_{\ell^{2}}$
with the matrix $\mathbf{A}$ and vector $\mathbf{F}$ describing the projection of the error $u-u_{i-1}$ onto the space spanned by the extended neural network activation functions. For ease of notation, we only state the linear system for the case when $L^{(i)}=1$ and note that extension to the case of an arbitrary depth network is straightforward (see Section 2.1):
[…]